### v1

@@ 8 Conclusion
With the release of different LLMs, especially those targeting different tasks in software engineering, practitioners and researchers have focused on properly using these models. Large Language Models such as ChatGPT and LLaMA are known for their potential to support users in their work, commonly treated as assistant tools. In this work, we investigate the potential of using LLMs as a replacement tool for Stack Overflow. We have conducted an empirical study assessing the reliability of the answers generated by these LLMs for existing questions in Stack Overflow while identifying possible challenges.
Overall, our results report a significant decline in user activity on Stack Overflow since the release of ChatGPT (questions, answers, and comments). However, for some topics, there is no statistical difference regarding the frequency of posted questions, showing that their communities keep active on Stack Overflow. Although ChatGPT outperforms LLaMA regarding the textual similarity of generated answers, LLaMA performs well, representing a good, strong, and free option to the general audience. While LLMs perform well on questions associated with general and specific problems on programming languages, they face some challenges when addressing frameworks and libraries questions. As a result, users might appeal to Stack Overflow when they do not get the expected support for their issues on LLMs.
These findings reinforce the power of LLMs to perform different tasks, bringing light to discussions regarding their impact. Replacing tools, like Stack Overflow for LLMs, represents a huge change that is not prudent to take for now. Currently, LLMs still have to further explore new features to provide generalizable, reliable, and valuable knowledge for their audience, while users are required to develop new skills to improve their experiences when adopting these models.
@@ 9 Acknowledgements
This work is funded by the following organizations and companies: Fonds de Recherche du Québec (FRQ), Natural Sciences and Engineering Research Council of Canada (NSERC), and Canadian Institute for Advanced Research (CIFAR). However, the findings and opinions expressed in this paper are those of the authors and do not necessarily represent or reflect those organizations/companies.
10 Data Availability
To promote open science and facilitate reproducibility, we make all our artifacts available to the community. This includes the datasets used in our experimentations, the source code for the scripts that were used, the results produced, and any other artifacts related to our study:
https://github.com/leusonmario/chat-stack

### v2

@@ 8 Conclusion
With the release of different LLMs, especially those targeting different tasks in software engineering, practitioners and researchers have investigated means for ensuring their proper usage. Large Language Models such as ChatGPT and LLaMA are known for their potential to support users in their work, commonly treated as assistant tools. In this work, we investigate the potential of using LLMs to address Stack Overflow questions. We have conducted an empirical study assessing the reliability of the answers generated by these LLMs for existing questions in Stack Overflow while identifying possible challenges.
Overall, our results report that answers generated by ChatGPT and LLaMA-2 show a high degree of textual similarity to the ones accepted in Stack Overflow. Although ChatGPT outperforms LLaMA regarding the textual similarity of generated answers, LLaMA performs well, representing a good, strong, and free option for the general audience. While LLMs perform well on questions associated with general and specific problems on programming languages, they face some challenges when addressing frameworks and libraries questions. As a result, users might appeal to Stack Overflow when they do not get the expected support for their issues on LLMs. Finally, we also observe a significant decline in user activity on Stack Overflow since the release of ChatGPT (questions, answers, and comments). Initially, for some domains, there is no statistically significant difference regarding the frequency of posted questions, showing that their communities keep active on Stack Overflow. However, after one year of the release, we observed a constant and significant decline for all domains evaluated.
These findings reinforce the capability of LLMs to perform different tasks, shedding light on discussions regarding their impact. Replacing tools, like Stack Overflow with LLMs, represents a huge change that is not prudent to take for now. Currently, LLMs still have to further explore new features to provide generalizable, reliable, and valuable knowledge for their audience, while users are required to develop new skills to improve their experiences when adopting these models.
@@ 9 Acknowledgements
We thank Arghavan Dakhel for her support in setting up the environment for this study. We also thank the anonymous reviewers for their valuable comments on improving an earlier version of this paper. This work is funded by the following organizations and companies: Fonds de Recherche du Quebec (FRQ), Natural Sciences and Engineering Research Council of Canada (NSERC), and the Canadian Institute for Advanced Research (CIFAR). However, the findings and opinions expressed in this paper are those of the authors and do not necessarily represent or reflect those organizations/companies.
Declaration
Conflicts of Interest: The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability: To promote open science and facilitate reproducibility, we make all our artifacts available to the community. This includes the datasets used in our experimentations, the source code for the scripts that were used, the results produced, and any other artifacts related to our study: https://github.com/leusonmario/chat-stack, https://doi.org/10.5281/zenodo.15086541.